OpenStudy (anonymous):

Use the quadratic formula to find any x-intercepts of the parabola. y=6x^2-5x+1

OpenStudy (anonymous):

Hey @wio i solved for zero after factoring and got x= 1/3, and x=1/2.. are those the x intercepts?

OpenStudy (anonymous):

You can do it.

OpenStudy (anonymous):

Plug them into the equation and see if you get 0.

OpenStudy (anonymous):

\[ 6\left(\frac 1 2\right)^2-5\left(\frac 1 2\right)+1 \]

OpenStudy (anonymous):

but.. isn't this right? http://answers.yahoo.com/question/index?qid=20090913014355AALEzOW

OpenStudy (anonymous):

lol. In precal in college and its review week. I forgot a lot over the summer but im pretty sure thats right...?

OpenStudy (anonymous):

I'm giving you a way to verify if they are correct.

OpenStudy (anonymous):

Otherwise you can just trust the quadratic formula worked.

OpenStudy (anonymous):

oh ok, yea. I would assume it works most of the time, lol. Because checking it would be some what annoying. lol